Suppose the demand and supply curves for a product is given by 𝑄 = 500 − 2𝑃 𝑄 = −100 + 3𝑃 a) Which is the supply curve and why? b) Graph the demand and supply curves. c) Compute the equilibrium price and quantity d) If the current price is 100, what is the quantity demanded and quantity supplied? How would you describe this situation and what would you expect to happen in this market? e) Suppose that the demand changes to 𝑄 = 600 − 2𝑃. Find the new equilibrium price and quantity and show this on your graph. 


1
Expert's answer
2022-03-11T08:40:34-0500

"Q=500-2P,"

"Q=-100+3P,"

"500-2P=-100+3P,"

"P=120,"

"Q=260."


"Q=600-2P,"

"Q=-100+3P,"

"600-2P=-100+3P,"

"P=140,"

"Q=320."